Find the polarizing angle of a dense flint glass having an index of 1.768
1
Expert's answer
2019-04-15T10:16:52-0400

In this case, Brewster's angle (also known as the polarization angle) is given by formula


"\theta=\arctan{\frac {n_2}{n_1}} (1)"

where n2=1.78, n1=1.00

Using (1) we got:


"\theta=60°"

Answer:

60°